# FOSP: Fine-tuning Offline Safe Policy through World Models

### v1

@@ 5.2 Main Results
Offline Pretrain
We show our offline training result in Figure 2. FOSP outperformed the similar safe algorithm SafeDreamer, excelling in decreasing costs and better balancing task performance with constraint avoidance. Furthermore, It has comparable or slightly lower performance than DreamerV3 while achieving nearly zero violations. Note that all methods struggle to get higher rewards in SafetyPointPush because the line-of-sight obstructions necessitate online exploration.
Online Fine-tuning
@@
The results of online fine-tuning are shown on Table 1. It proves that our algorithm not only performs well offline but can also be fine-tuned to achieve better performance. Meanwhile, FOSP further reduces the costs during the online fine-tuning. We also compare our method with the OSRP version of SafeDreamer which can only be trained online. We notice that the SafeDreamer training exclusively online outperforms the offline-to-online fine-tuning approach. This may be because direct fine-tuning is affected by distribution shifts, making it difficult to improve performance in limited steps. So FOSP is necessary for us to bridge the safety of offline and online while maintaining its performance. (For more detail see Appendix.F)
Ablation Studies
@@
In this section, we compare the effects of different components of the model on its performance in Figure 3. Specifically, we provide the following ablations: (i) learning value function without in-sample optimization (Eq.6) in Section 4.1; (ii) directly fine-tuning the model without model-based policy expansion (Eq.9) in Section 4.2; (iii) update safe actor without reachability estimation function in Section 4.3; (iv) design the penalty without the Augmented Lagrangian and use the feasible value function[52]. The results show that each component of our method individually contributes to ensuring offline training performance, further improvement during online fine-tuning, and maintaining near-zero cost constraint violations.
Table 1: Online fine-tuning results. Reward return and cost return of methods in offline-to-online fine-tuning (1M steps for offline and 0.5M steps for online fine-tuning). And results of online algorithms only trained online (1.5M steps). Means of three seeds.
Offline-to-Online Online-only
DreamerV3 SafeDreamer FOSP(Ours) SafeDreamer(+planning)
Tasks Reward Cost Reward Cost Reward Cost Reward Cost
PointGoal1 17.4$\rightarrow$25.8 82.3$\rightarrow$85.1 12.7$\rightarrow$19.2 1.5$\rightarrow$0.02 18.7$\rightarrow$21.5 1.4$\rightarrow$0.2 20.1 0.6
PointButton1 15.1$\rightarrow$20.8 167.7$\rightarrow$157.9 8.9$\rightarrow$10.7 10.2$\rightarrow$4.5 14.7$\rightarrow$18.1 9.5$\rightarrow$2.1 12.4 5.5
PointPush1 2.3$\rightarrow$14.6 30.9$\rightarrow$26.3 1.3$\rightarrow$10.3 17.8$\rightarrow$0.12 4.0$\rightarrow$13.2 18.1$\rightarrow$0.1 8.1 0.7
PointGoal2 9.5$\rightarrow$18.9 367.1$\rightarrow$290.2 3.6$\rightarrow$12.7 8.6$\rightarrow$0.3 8.1$\rightarrow$13.5 7.5$\rightarrow$0.23 13.0 1.7
CarGoal2 16.7$\rightarrow$24.1 231.0$\rightarrow$287.9 6.9$\rightarrow$9.8 0.7$\rightarrow$0.09 10.1$\rightarrow$14.5 1.6$\rightarrow$0.07 11.2 0.34
Average 20.8 171.1 12.5 1.01 16.6 0.54 12.9 1.77
5.3 Unseen Safe Tasks
In the real-world environment, robots will incur many safety constraints while performing various tasks. Our key challenge is how to safely train the robots to adapt to unseen constraints by safe learning-based methods. Following this insight, we introduce obstacles and targets of different shapes and sizes in Reach tasks, where the robot has only encountered a limited number of these during offline training. Our goal is to make the robot use a few trials to learn planning strategies in environments with new obstacles as safely as possible. Thus, we design three real-world transfer tasks (see details in Appendix.C) and show their results in Table 2.
@@ 5.3 Unseen Safe Tasks
Table 2: Real-world unseen tasks. We record the success rate (SR, %) and the constraint violation rate (CV, %) over 20 tests in three tasks while it will be labeled as a violation if it collides with an obstacle. The robot has fine-tuned 40 steps. See Appendix.C for more details.
@@
1 40 50 45 40 50 30
2 30 50 35 45 35 40
3 20 60 40 40 50 30
@@ 6 Conclusion
We introduce FOSP, a novel method that integrates model-based reinforcement learning with offline-to-online fine-tuning to enhance safety in robotic tasks. FOSP addresses the challenges of balancing performance and safety during both offline and online phases using the reachability estimation function to unify different constraints and offline safe policies to guide online exploration. Our experimental results demonstrate that FOSP achieves robust performance in various safe planning tasks, especially those with high-dimensional vision inputs. Furthermore, we evaluate FOSP on unseen safety-critical tasks in dynamic real-world environments highlighting its practical value in safe few-shot fine-tuning. FOSP is the first approach to extend MBRL into safe offline-to-online training, showing promising results in the Safety-Gymnasium benchmark. Overall, FOSP shows a significant step forward in ensuring safe and efficient decision-making in embodied intelligence.
@@ A.1 Derivation of the Final Optimization Problem
This section analyzes the derivation from Eq.(21) to Eq.(13) by using the Augmented Lagrangian and advantage function. First, we introduce Proposition 1.
@@
The optimization objective of Eq.(11) is the necessary condition of $\max_{\pi}\operatorname*{\mathbb{E}}_{{\bm{s}}}[A^{r}({\bm{s}},{\bm{a}})\cdot\mathds{1}_{s\in\mathcal{S}_{f}}]$.
Proof. We start with the advantage function estimation in [53] and [23]. Suppose that $d_{\pi}({\bm{s}})=\sum_{t=0}^{\infty}\gamma^{t}p({\bm{s}}_{t}={\bm{s}}|\pi)$ is discounted state distribution with $\pi$, and $p({\bm{s}}_{t}={\bm{s}}|\pi)$ is the probability of the state ${\bm{s}}$ guided by $\pi$ for $t$ steps. The discounted sum of advantage expectation under $d_{\pi}({\bm{s}})$ can be represented as follows:
$\operatorname*{\mathbb{E}}_{{\bm{s}}_{1:\infty}\sim d_{\pi}({\bm{s}})}\left[\sum_{t=0}^{\infty}\gamma^{t}A_{\mu}({\bm{s}}_{t},{\bm{a}}_{t})\right]$
$=\operatorname*{\mathbb{E}}_{{\bm{s}}_{1:\infty}\sim d_{\pi}({\bm{s}})}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(r({\bm{s}}_{t},{\bm{a}}_{t})+\gamma V_{\mu}({\bm{s}}_{t+1})-V_{\mu}({\bm{s}}_{t})\right)\right]$
$=\operatorname*{\mathbb{E}}_{{\bm{s}}_{1:\infty}\sim d_{\pi}({\bm{s}})}\left[-V_{\mu}({\bm{s}}_{0})+\sum_{t=0}^{\infty}\gamma^{t}r({\bm{s}}_{t},{\bm{a}}_{t})\right]$
$=-\operatorname*{\mathbb{E}}_{{\bm{s}}_{0}\sim d_{\pi}({\bm{s}}_{0})}\left[V_{\mu}({\bm{s}}_{0})\right]+\operatorname*{\mathbb{E}}_{{\bm{s}}_{1:\infty}\sim d_{\pi}({\bm{s}})}\left[\sum_{t=0}^{\infty}\gamma^{t}r({\bm{s}}_{t},{\bm{a}}_{t})\right]$
@@
where $\mu$ is the behavior policy used to sample data. So we get the objective representation $J(\pi)=\operatorname*{\mathbb{E}}_{{\bm{s}}\sim d_{\pi}({\bm{s}})}\Big{[}\operatorname*{\mathbb{E}}_{{\bm{a}}\sim\pi({\bm{a}}|{\bm{s}})}[A_{\mu}({\bm{s}}_{t},{\bm{a}}_{t})]\Big{]}+J(\mu)$. Then, we have the following derivation:
$\max_{\pi}\operatorname*{\mathbb{E}}_{{\bm{s}}\sim d_{\pi}({\bm{s}})}[A^{r}_{\pi}({\bm{s}},{\bm{a}})]$ $\Rightarrow\max_{\pi}\operatorname*{\mathbb{E}}_{{\bm{s}}\sim d_{\pi}({\bm{s}})}\Big{[}\operatorname*{\mathbb{E}}_{{\bm{a}}\sim\pi({\bm{a}}|{\bm{s}})}[A_{\mu}^{r}({\bm{s}}_{t},{\bm{a}}_{t})\Big{]}$
$=\max_{\pi}\operatorname*{\mathbb{E}}_{{\bm{s}}\sim d_{\pi}({\bm{s}})}\Big{[}\operatorname*{\mathbb{E}}_{{\bm{a}}\sim\pi({\bm{a}}|{\bm{s}})}[A_{\mu}^{r}({\bm{s}}_{t},{\bm{a}}_{t})\Big{]}+J(\mu)$
@@
$=\max_{\pi}\;\operatorname*{\mathbb{E}}[V_{\pi}^{r}({\bm{s}})].$
The last equation holds by the definition of $V_{\pi}^{r}({\bm{s}})$. $\square$
Due to the symmetry of $\max_{\pi}\operatorname*{\mathbb{E}}[V_{\pi}^{r}({\bm{s}})]$ and $\max_{\pi}\operatorname*{\mathbb{E}}[V_{\pi}^{c}({\bm{s}})]$, the same proof also applies to $\max_{\pi}\operatorname*{\mathbb{E}}[V_{\pi}^{c}({\bm{s}})]$. So we have proposition 2.
@@
The optimization objective of Eq.(12) is the necessary condition of $\max_{\pi}\operatorname*{\mathbb{E}}_{{\bm{s}}}[-A^{c}({\bm{s}},{\bm{a}})\cdot\mathds{1}_{s\notin\mathcal{S}_{f}}]$.
The problem converts to:
@@
However, it is hard for us to get a close form of $\pi_{\psi}$ in this problem. Therefore, we will first disregard the constraints of the behavior policy and address the hard safety constraints. The Augmented Lagrangian is used to deal with the feasible part. Thus, we use the following relaxation:
$\max_{\pi_{\psi}}\min_{\lambda\geq 0}\left[A^{r}({\bm{s}},{\bm{a}})\mathds{1}_{s\in\mathcal{S}_{f}}-\lambda V_{\varphi}^{c}({\bm{s}})+\frac{1}{\mu_{k}}(\lambda-\lambda_{k})^{2}\right],$ (20)
where $\mu_{k}$ is a non-decreasing penalty term corresponding to gradient step $k$. By smoothing the left-hand side term in Eq.(18), the last term ensures that $\lambda$ stays near its previous estimate. Differentiating Eq.(18) with respect to $\lambda_{p}^{k}$ leads to the following update rule for the Lagrange multiplier:
$\lambda_{p}^{k+1}=\begin{cases}\lambda_{p}^{k}+\mu^{k}V_{\varphi}^{c}&\text{ if }\lambda_{p}^{k}+\mu^{k}V_{\varphi}^{c}\geq 0,\\ 0&\text{ otherwise }.\end{cases}$
The objective becomes the following form:
$J(\pi,\lambda_{k},\mu_{k})=A^{r}({\bm{s}},{\bm{a}})\mathds{1}_{s\in\mathcal{S}_{f}}-\Psi\left(V_{\varphi}^{c},\lambda_{p}^{k},\mu^{k}\right),$
$\Psi\left(V_{\varphi}^{c},\lambda_{p}^{k},\mu^{k}\right)=\begin{cases}\lambda_{p}^{k}V_{\varphi}^{c}+\frac{\mu^{k}}{2}(V_{\varphi}^{c})^{2}&\text{ if }\lambda_{p}^{k}+\mu^{k}V_{\varphi}^{c}\geq 0,\\ -\frac{\left(\lambda_{p}^{k}\right)^{2}}{2\mu^{k}}&\text{ otherwise }.\end{cases}$
Then, we utilize the REF function to connect the two parts of the problem. The REF function is defined as follows:
Definition 1.
The reachability estimation function (REF) $u:\mathcal{S}\rightarrow[0,1]$ is:
$u^{\pi}({\bm{s}}):=\operatorname*{\mathbb{E}}_{\tau\sim\pi}[\max_{{\bm{s}}_{t}\in\tau}\mathds{1}_{({\bm{s}}_{t}|{\bm{s}}_{0},\pi)\in\mathcal{S}_{f}}]$
And the problem becomes:
$\max_{\pi_{\psi}}\;\operatorname*{\mathbb{E}}_{{\bm{s}}}\Big{[}V_{\varphi}^{r}({\bm{s}})\cdot u^{\pi}_{\psi}-V_{\varphi}^{c}({\bm{s}})\cdot(1-u^{\pi}_{\psi})\Big{]},\;\;\text{s.t.}\;\;V_{\varphi}^{c}({\bm{s}})=0,\;\;D(\pi_{\psi}||\pi_{b})\leq\epsilon.$ (21)
We introduce a proposition from [10] to solve the soft constraints from offline training to provide a solution of Eq.(13).

### v2

@@ 5.2 Main Results
Figure 5: Simulation generalization task. We compared the performance of models pretrained 1M steps on SafetyPointGoal1 using FOSP and SafeDreamer, evaluating their fine-tuning and generalization results after 0.5M steps on SafetyPointGoal2.
@@
We show our main offline pretrain and online fine-tuning results in Table 1, 2 to answer the question (1). During offline pretraining, FOSP outperformed the similar safe algorithms Recover RL and SafeDreamer, excelling in decreasing costs and better balancing task performance with constraint avoidance. Furthermore, it has comparable or slightly lower performance than DreamerV3 with nearly zero violations. (Appendix.H) The learning curves of online fine-tuning are shown in Figure 3. It proves that our algorithm can also be fine-tuned to achieve better performance. Meanwhile, FOSP further reduces the costs during the online fine-tuning. We also compare our method with the online planning version of SafeDreamer (OSRP) (Huang et al., 2023) which can only be trained online. We notice that the SafeDreamer training exclusively online outperforms the offline-to-online approach but has more constraint violations. This may be because the safety constraints learned during offline pretraining provide guidance for online fine-tuning, making the model more sensitive to safety constraints. Taking this advantage, FOSP is able to bridge offline and online safety while maintaining its performance.
5.3 Ablation Studies
Module Ablation
@@ 5.3 Ablation Studies
In this section, we compare the effects of different components of the model on its performance in Figure 4 to answer the question (2). Specifically, we provide the following ablations: (i) learning value function without in-sample optimization (equation 7) in Section 4.1; (ii) directly fine-tuning the model without safe policy expansion (equation 21) in Section 4.3; (iii) update safe actor without reachability estimation function in Section 4.2; (iv) design the penalty without the Augmented Lagrangian and use the feasible value function (Fisac et al., 2019). The results show that each component of our method individually (i) improves offline performance, (ii) stabilizes the transition phase from offline to online, (iii) enhances overall performance while maintaining safety constraints, and (iv) ensures compliance with safety constraints.
@@
Table 3: Real-world unseen tasks. We record the success rate (SR, %) and the constraint violation rate (CV, %) over 20 tests in three tasks while it will be labeled as a violation if it collides with an obstacle. The robot has fine-tuned 40 gradient steps. See Appendix.E for more details.
@@
1 40 50 45 40 65 20
2 30 50 35 45 60 30
3 20 60 40 40 50 20
Figure 6: Dataset ablation studies. We evaluate ablations on SafetyPointGoal2 with five seeds. Due to the large offline costs affecting the visual presentation, we reduced them by 10 times in the left figures. The x label in the left image represents safe data : unsafe data. 1M steps for offline and 0.5M steps for online fine-tuning.
Dataset Ablation
We also seek to investigate how the coverage and size of offline datasets affect the algorithm’s performance to answer the question (3). For the dataset coverage, it is well established that higher rewards in the dataset lead to better performance (Yu et al., 2020). Therefore, we focus on examining the impact of cost by adjusting the proportion of safe and unsafe data. The results in Figure 6 demonstrate that safe data tends to guide the final policy towards being overly aggressive, achieving high rewards but frequently violating constraints. On the other hand, unsafe data causes the policy to become overly conservative, hindering exploration to avoid constraint violations. This behavior arises due to errors in the dynamics model in model-based methods. When the dataset is imbalanced, the dynamics model trained through supervised learning struggles to accurately predict the costs associated with each state-action pair in the real environment. As a result, the actor-critic, which depends heavily on data generated by the dynamics model, fails to predict future state costs accurately, resulting in a suboptimal policy. To examine the impact of dataset size, we scaled the standard dataset up and down proportionally. The final results indicate that while increasing the dataset size leads to the learning of better and safer policies. Furthermore, the gains during the online fine-tuning stage are relatively modest.
5.4 Safe Generalization Tasks
In this section, we show how our method solves different tasks with unseen safety regions in the offline dataset and answer the question (4). We first evaluate it in the simulation, offline pretraining the agent using data collected in the SafetyPointGoal1, then place it in the SafetyPointGoal2, which has more types and a greater number of unsafe regions, for online fine-tuning. Although these tasks are the same type, the limitations of the previous data still pose significant challenges to the effectiveness and safety of fine-tuning. The experimental results in Figure 5 demonstrate the advantages of FOSP in these safety generalization tasks, showing better performance and safer fine-tuning compared to SafeDreamer.
In the real-world environment, robots will encounter many safety constraints while performing various tasks. Our key challenge is how to safely train the robots to adapt to unseen constraints by safe learning-based methods. Following this insight, we introduce obstacles and targets of different shapes and sizes in Reach tasks, where the robot will only encounter a limited number of these during offline pretraining. We first train the agent for 0.5 million steps using a precollected dataset, then deploy the trained agent to the real-world environment to execute tasks and fine-tune it. During online fine-tuning, each time the robot finishes a trajectory, we add it to the replay buffer and update model parameters. Our goal is to make the robot use a few trials to learn planning policies in environments with new obstacles as safely as possible. Thus, we design three real-world transfer tasks and show their results in Table 3. For more details see Appendix E.
@@ 6 Conclusion
We introduce FOSP, a novel method that integrates model-based reinforcement learning with offline-to-online fine-tuning to enhance safety in robotic tasks. FOSP addresses the challenges of balancing performance and safety during both offline and online phases using the reachability estimation function to unify different constraints and offline safe policies to guide online exploration. Our experimental results demonstrate that FOSP achieves robust performance in various safe planning tasks with vision inputs. Furthermore, we evaluate FOSP on unseen safety-critical tasks in both simulation and dynamic real-world environments, highlighting its practical value in safe few-shot fine-tuning. Overall, FOSP is the first approach to extend MBRL into a safe offline-to-online framework to solve the safe generalization problem, showing promising results in the Safety-Gymnasium benchmark.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (No.62103225), the Natural Science Foundation of Guangdong Province (No.2024A1515010003) and the Natural Science Foundation of Shenzhen (No.JCYJ20230807111604008, No. JCYJ20240813112007010).
@@ A.1 Derivation of the Final Optimization Problem
This section analyzes the derivation from equation 13, equation 14 to equation 16 by using the Augmented Lagrangian and advantage function. First, we introduce Proposition 1, which refers to Zheng et al. (2023) (Theorem 1), Peng et al. (2019) (Appendix A).
@@
The optimization objective of equation 13 is the necessary condition of $\max_{\pi}\mathbb{E}_{{\bm{s}}}[A^{r}({\bm{s}},{\bm{a}})\cdot\mathds{1}\{s\in\mathcal{S}_{f}\}]$.
Proof. We start with the advantage function estimation in Kakade & Langford (2002) and Peng et al. (2019). Suppose that $d_{\pi}({\bm{s}})=\sum_{t=0}^{T}\gamma^{t}p({\bm{s}}_{t}={\bm{s}}|\pi)$ is discounted state distribution with $\pi$, and $p({\bm{s}}_{t}={\bm{s}}|\pi)$ is the probability of the state ${\bm{s}}$ guided by $\pi$ for $t$ steps. The discounted sum of advantage expectation under $d_{\pi}({\bm{s}})$ can be represented as follows:
$\mathbb{E}_{{\bm{s}}_{0:T}\sim d_{\pi}({\bm{s}})}\left[\sum_{t=0}^{T}\gamma^{t}A_{\mu}({\bm{s}}_{t},{\bm{a}}_{t})\right]$
$=\mathbb{E}_{{\bm{s}}_{0:T}\sim d_{\pi}({\bm{s}})}\left[\sum_{t=0}^{T}\gamma^{t}\left(r({\bm{s}}_{t},{\bm{a}}_{t})+\gamma V_{\mu}({\bm{s}}_{t+1})-V_{\mu}({\bm{s}}_{t})\right)\right]$
$=\mathbb{E}_{{\bm{s}}_{0:T}\sim d_{\pi}({\bm{s}})}\left[-V_{\mu}({\bm{s}}_{0})+\sum_{t=0}^{T}\gamma^{t}r({\bm{s}}_{t},{\bm{a}}_{t})\right]$
$=-\mathbb{E}_{{\bm{s}}_{0}\sim d_{\pi}({\bm{s}}_{0})}\left[V_{\mu}({\bm{s}}_{0})\right]+\mathbb{E}_{{\bm{s}}_{1:T}\sim d_{\pi}({\bm{s}})}\left[\sum_{t=0}^{T}\gamma^{t}r({\bm{s}}_{t},{\bm{a}}_{t})\right]$
@@
where $\mu$ is the behavior policy used to sample data. So we get the objective representation $J(\pi)=\mathbb{E}_{{\bm{s}}\sim d_{\pi}({\bm{s}})}\Big{[}\mathbb{E}_{{\bm{a}}\sim\pi({\bm{a}}|{\bm{s}})}[A_{\mu}({\bm{s}}_{t},{\bm{a}}_{t})]\Big{]}+J(\mu)$. Then, we have the following derivation:
$\max_{\pi}\mathbb{E}_{{\bm{s}}\sim d_{\pi}({\bm{s}})}[A^{r}_{\pi}({\bm{s}}_{t},{\bm{a}}_{t})]$ $\Rightarrow\max_{\pi}\mathbb{E}_{{\bm{s}}\sim d_{\pi}({\bm{s}})}\Big{[}\mathbb{E}_{{\bm{a}}\sim\pi({\bm{a}}|{\bm{s}})}[A_{\mu}^{r}({\bm{s}}_{t},{\bm{a}}_{t})]\Big{]}$
$=\max_{\pi}\mathbb{E}_{{\bm{s}}\sim d_{\pi}({\bm{s}})}\Big{[}\mathbb{E}_{{\bm{a}}\sim\pi({\bm{a}}|{\bm{s}})}[A_{\mu}^{r}({\bm{s}}_{t},{\bm{a}}_{t})]\Big{]}+J(\mu)$
@@
$=\max_{\pi}\;\mathbb{E}[V_{\pi}^{r}({\bm{s}}_{t})].$
The last equation holds by the definition of $V_{\pi}^{r}({\bm{s}}_{t})$. $\square$
Due to the symmetry of $\max_{\pi}\mathbb{E}[V_{\pi}^{r}({\bm{s}})]$ and $\max_{\pi}\mathbb{E}[V_{\pi}^{c}({\bm{s}})]$, the same proof extends to $\max_{\pi}\mathbb{E}[V_{\pi}^{c}({\bm{s}})]$ as well. So we have proposition 2.
@@
The optimization objective of equation 14 is the necessary condition of $\max_{\pi}\mathbb{E}_{{\bm{s}}}[-A^{c}({\bm{s}},{\bm{a}})\cdot\mathds{1}\{s\notin\mathcal{S}_{f}\}]$.
The problem (equation 13, equation 14) converts to:
@@
However, it is hard for us to get a close form of $\pi_{\psi}$. Therefore, we begin by disregarding the constraints of the behavior policy and focus on addressing the hard safety constraints. The Augmented Lagrangian is used to deal with the feasible part (As et al., 2022). We use the following relaxation:
$\max_{\pi_{\psi}}\min_{\lambda_{p}\geq 0}\left[A^{r}({\bm{s}},{\bm{a}})\mathds{1}\{{\bm{s}}\in\mathcal{S}_{f}\}-\lambda_{p}V_{\varphi}^{c}({\bm{s}})+\frac{1}{\mu^{k}}(\lambda_{p}-\lambda_{p}^{k})^{2}\right],$ (26)
where $\lambda_{p}$ is a Lagrange multiplier and $\mu^{k}$ is a non-decreasing penalty term corresponding to gradient step $k$. By smoothing the left-hand side term in equation 24, the last term ensures that $\lambda_{p}$ stays near its previous estimate. Differentiating equation 24 with respect to $\lambda_{p}^{k}$ leads to the following update rule for the Lagrange multiplier:
$\lambda_{p}^{k+1}=\begin{cases}\lambda_{p}^{k}+\frac{\mu^{k}}{2}V_{\varphi}^{c}&\text{ if }\lambda_{p}^{k}+\frac{\mu^{k}}{2}V_{\varphi}^{c}\geq 0,\\ 0&\text{ otherwise }.\end{cases}$ (27)
where the $\lambda_{p}^{k+1}$ only update when $\pi_{\psi}$ satisfies the constraints. The feasible objective becomes the following form:
$J(\pi_{\psi},\lambda_{p}^{k},\mu^{k})=A^{r}({\bm{s}},{\bm{a}})\mathds{1}\{{\bm{s}}\in\mathcal{S}_{f}\}-\Phi\left(V_{\varphi}^{c},\lambda_{p}^{k},\mu^{k}\right),$ (28)
$\Phi\left(V_{\varphi}^{c},\lambda_{p}^{k},\mu^{k}\right)=\begin{cases}\lambda_{p}^{k}V_{\varphi}^{c}+\frac{\mu^{k}}{4}(V_{\varphi}^{c})^{2}&\text{ if }\lambda_{p}^{k}+\frac{\mu^{k}}{2}V_{\varphi}^{c}\geq 0,\\ -\frac{\left(\lambda_{p}^{k}\right)^{2}}{\mu^{k}}&\text{ otherwise }.\end{cases}$ (29)
Then, we utilize the reachability estimation function to connect the two parts of the problem like 15. So the final problem becomes equation 16.
A.2 Extraction of the Optimal Policy
We introduce a proposition from Nair et al. (2020) to solve the soft constraints from offline training to provide a solution of equation 16.